OpenStudy (anonymous):

You invest $200 at 4% compounded annually. How many years would it take for the $200 to double in value?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

help help help !!! @surjithayer

OpenStudy (anonymous):

help me get the answer

OpenStudy (anonymous):

\[400=200\left( 1+\frac{ 4 }{ 100 } \right)^{n}\] \[\frac{ 400 }{ 200 }=\left( \frac{ 26 }{ 25 } \right)^{n},2=\left( \frac{ 26 }{ 25 }\right)^{n}\] taking log of both sides and solve.

OpenStudy (anonymous):

how do I take log from both side and solve ? Im really confused .. can you just show me the whole step?

OpenStudy (anonymous):

log1.04( 200-1.04) ???

OpenStudy (anonymous):

\[\log \left( \frac{ 26 }{ 25 } \right)^{n}=\log2,n \left( \log26-\log25 \right)=\log2,n=\frac{ \log 2 }{ \log26-\log25 }\]

OpenStudy (anonymous):

the answer is ..17.67???

OpenStudy (anonymous):

now you can solve by taking log

OpenStudy (anonymous):

answer ^^^^

OpenStudy (anonymous):

17.67?

OpenStudy (anonymous):

i have no log table or calculator. you can solve.

OpenStudy (anonymous):

try this http://www.wolframalpha.com/input/?i=log+fit+calculator

OpenStudy (anonymous):

I want you to check if it's right

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

is it 18 years ?????@surjithayer

OpenStudy (anonymous):

\[n=\frac{ .6931 }{ 3.2580-3.2189}=17.67\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw